OpenStudy (anonymous):

Horizontal translation of a quartic function by 30 y=-x^3+138x

OpenStudy (anonymous):

I need to figure how how I can translate this equation to the right by 30

OpenStudy (anonymous):

I know this looks pretty rudimentary but trust me I don't know how

zepdrix (zepdrix):

Hey Robert :) In general, to make an horizontal adjustment, we would compensate by doing the opposite to our x. So here is an example with a simple quadratic.\[\large\rm f(x)=(x)^2\]To shift it 30 to the right (positive), ill replace x with x-30,\[\large\rm g(x)=(x-30)^2\]This new function g(x) is the same shape as f(x) but the whole thing has been shifted to the right 30 units.

zepdrix (zepdrix):

So we have a cubic function and we want to apply the same idea, (not a quartic silly :3 ) \[\large\rm y_1=-\color{orangered}{x}^3+138\color{orangered}{x}\]\[\large\rm y_2=-\color{orangered}{(x-30)}^3+138\color{orangered}{(x-30)}\]And then maybe you'll have to expand some brackets out if you want to get it into standard form.

OpenStudy (anonymous):

Oh I see it now. I have to apply the same shift to all the variables.

OpenStudy (anonymous):

Thanks so much @zepdrix Are you taking university math?

zepdrix (zepdrix):

Yesss :) Learning lots of fun stuff!!